OpenStudy (mymomlovesme):

Matthew is 5 years older than twice the age of his sister. If Matthew is 13, how old is his sister?

OpenStudy (mymomlovesme):

Help!! Will medal and fan! Multiple choice: a. 3 b. 4 c. 9 d. 31

OpenStudy (calebb8133):

This is 5th grade math come on!!!!!!!!

OpenStudy (brainzonly):

love these! Well, first make an equation

OpenStudy (calebb8133):

this is to easy for open study

OpenStudy (mymomlovesme):

Uh..

OpenStudy (mymomlovesme):

I know XD My brother asked it.

OpenStudy (brainzonly):

shut up caleb stop being mean! have you ever read the code of conduct!?? 2x+5 5 years older than TWICE the age of his sister, x

OpenStudy (calebb8133):

then u figure it out yourself if your brother gave it to U

OpenStudy (brainzonly):

5+2x=13 solve that

OpenStudy (mymomlovesme):

x= 4

OpenStudy (brainzonly):

so B XD now put 4 in for x does that make a true statement? 5+ 2x = 13 5 + 2(4) = 13 5+ 8 = 13 13=13 yes.. x = 4

OpenStudy (mymomlovesme):

Okay never mind but thanks.

OpenStudy (mymomlovesme):

Oh I didn't see that reply after x=4 lmao

OpenStudy (mymomlovesme):

Thanks :/

OpenStudy (brainzonly):

np :) have a great day

OpenStudy (mymomlovesme):

You too ;)